OpenStudy (anonymous):

lim x approaches 0 (sinx-tanx)

OpenStudy (anonymous):

Sine and Tangent are both continuous at a neighbourhood of 0 therefore we can just plug in the value 0: \[\lim_{x\to0}(\sin x-\tan x)=(\sin 0-\tan 0)=0-0=0\]